OpenStudy (anonymous):

Solve the DiffEQ (y^2+xy^2)y'=1

OpenStudy (jamesj):

This equation is separable. Separate and integrate.

OpenStudy (anonymous):

Tried to split don't know how to deal with y^2dy/dx+xy^2dy/dx

OpenStudy (anonymous):

unless I can change the equation to y^2+xy^2= dx/dy...

OpenStudy (jamesj):

Think for a minute. You want to write y^2 + xy^2 as the product of a function in y and a function in x. How can you factorize this expression in order to do so?

OpenStudy (anonymous):

2y^2dy=dx/x?

OpenStudy (jamesj):

No, that simply can't be right by basic algebra. Factorize y^2 + xy^2 for me.

OpenStudy (anonymous):

(x+1)y^2?

OpenStudy (jamesj):

Yes ... hence .... now separating, you have ... what?

OpenStudy (anonymous):

What do I do with the y'?

OpenStudy (jamesj):

(x+1)y^2 dy/dx = 1. Therefore y^2 dy = dx / (x + 1)

OpenStudy (anonymous):

Oh I see.